OpenStudy (anonymous):

Why can't x1 be equal to x2?

Directrix (directrix):

>>Why can't x1 be equal to x2? Please post the entire question.

OpenStudy (nechirwan):

\[X ^{1} =X ^{2}\]

Directrix (directrix):

If the 2 and 2 are exponents and not subscripts, then there are values of x for which the two expressions are equal.

OpenStudy (mathmath333):

let \[x=1\] then\[x^1=1^1=1\\x^2=1^2=1\]

OpenStudy (mathmath333):

in this case it is equal

OpenStudy (mathmath333):

and in the case \[x=0\]

OpenStudy (jhannybean):

x is one representation for all number values, so the reason they cannot equal eachother is because x is not defined as one set number for both functions.